Federal Reserve Bank of Minneapolis Research Department Staff Report 328

Revised November 2006

JOURNAL FIGURES: Business Cycle Accounting^{†}

V.V. Chari University of Minnesota and Federal Reserve Bank of Minneapolis

Patrick J. Kehoe Federal Reserve Bank of Minneapolis and National Bureau of Economic Research

Ellen R. McGrattan Federal Reserve Bank of Minneapolis and University of Minnesota

† The views expressed herein are those of the author and not necessarily those of the Federal Reserve Bank of Minneapolis or the Federal Reserve System.